# When AIOps Become “AI Oops”:

### v1

@@ Abstract
In this work, we perform the first security analysis of AIOps solutions, showing that, once again, AI-driven automation comes with a profound security cost. We demonstrate that adversaries can manipulate system telemetry to mislead AIOps agents into taking actions that compromise the integrity of the infrastructure they manage. We introduce techniques to reliably inject telemetry data using error-inducing requests that influence agent behavior through a form of adversarial input we call adversarial reward-hacking; plausible but incorrect system error interpretations that steer the agent’s decision-making. Our attack methodology, AIOpsDoom, is fully automated, combining reconnaissance, fuzzing, and LLM-driven adversarial input generation–and operates without any prior knowledge of the target system.
@@ 1 Introduction
The increasing sophistication of software systems has led organizations to integrate AI deeply into IT operations, a paradigm widely known as AIOps, or AI for IT Operations [12]. At its core, AIOps leverages machine learning methods to automate tasks traditionally performed by human operators, including anomaly detection, incident diagnosis, and automated remediation [56]. The new wave of AIOps incorporates autonomous agents built upon large language models (LLMs), which dynamically interact with systems to diagnose issues and execute corrective actions. This rapid adoption of AI-driven automation promises significant benefits for IT Operations, notably reduced response times, increased efficiency, and lower operational costs; a vision shared by both academia [11, 43, 53, 10, 52, 28, 50, 51, 47, 49, 45] and industry [13, 44, 19, 34, 42, 20, 32, 40, 57, 24, 18].
@@
(In)security Through Automation However, existing research [11, 43, 53, 10, 52, 28, 50, 51, 47, 49, 45] has generally overlooked the security impact that this automation introduces. AIOps agents rely heavily on the telemetry data, which they consume to make decisions. So far in this area, the telemetry data has been assumed to be faithful and trustworthy. In this work, we challenge the above (false) assumption. We demonstrate that an adversary can manipulate telemetry data to indirectly influence and control the behavior of the AIOps agent.
Specifically, we present the very first security assessment of AIOps. We show that even in the most realistic threat models in which the attacker knows nothing about the system or the agent, they can pollute telemetry to bias AIOps agents into executing harmful remediations and compromise production systems. The consequences can be severe, enabling attackers to subvert secure systems through the very automation designed to protect them.
At a more technical level, the proposed attack methodology, which we call AIOpsDoom, is based on a collection of known and novel components that collectively break state-of-the-art open-source AIOps solutions. Using classical reconnaissance techniques, the attacker collects information about the system under attacks solely through its publicly available interfaces. Our methodology in AIOpsDoom identifies which entry points are meaningful for injecting an adversarially crafted payloads, so as to end up in the stored telemetry data that the AI agent will later process. The goal is to design payloads that will trigger the AI agent to bring the system to an insecure state voluntarily, thereby introducing a vulnerability through this automation pipeline.
Differences from Prompt Injection Approaches. Prompt injection refers to a class of attacks in which a maliciously crafted input manipulates the LLM or agent, causing it to completely override its intended task and act arbitrarily. While modern LLMs have developed improved resilience against traditional prompt injection techniques, making such attacks harder to execute in practice, our work explores a distinct strategy. We introduce a more subtle and tailored form of manipulation–one that preserves the apparent legitimacy of the agent’s task while covertly influencing its behavior. We draw inspiration from the concept of reward hacking [7, 16], where an agent exploits underspecified objectives or environment to maximize rewards through low-effort yet reward-generating behaviors. Our approach designs payloads embedded in telemetry data that present plausible interpretations of system errors, a technique we call adversarial reward-hacking. We evaluate our attack techniques against state-of-the-art models such as GPT-4o and GPT-4.1, and show that they can evade sophisticated prompt-defense solutions, including Microsoft’s PromptShields and Meta’s PromptGuard-2.
@@
  • We present the first security assessment of Agentic AI in the context of IT Operations, known as AIOps. Our end-to-end attack methodology integrates customized adversarial techniques, drawing inspiration from well-established attack vectors, software testing strategies, and reconnaissance principles.
  • At the core of our attack methodology lies a technique called AIOpsDoom, which serves a dual purpose. First, it enumerates all potential client interactions where manual data entry is permitted, a task handled by the Crawler component of AIOpsDoom. Second, the list of identified interactions that accept user data is passed to the component of AIOpsDoom called Fuzzer, which identifies which of them can cause an error, an action that can write the client data directly to the system’s telemetry data.
@@
  • We propose AIOpsShield, a defense mechanism tailored for AIOps that sanitizes telemetry data by leveraging its structured nature and the minimal influence of user-generated content in legitimate incident management tasks. Empirical evaluations on established benchmarks show that AIOpsShield effectively blocks telemetry-based adversarial attacks without degrading agent performance.
@@ 2.1 Telemetry and Observability
Observability tools collect, analyze, and correlate telemetry data to provide insights into the internal states of IT systems. Any modern and sufficiently complex IT architecture today relies on some form of observability stack (a combination of multiple observability tools) to maintain a comprehensive view of the system, facilitate incident response, and enable effective root cause analysis.
Collected data falls into three main categories:
@@
Numerous vendors offer observability solutions, either as on-premise products or as cloud-based observability platforms provided as services. While different products may feature unique functionalities, core capabilities typically remain consistent across these solutions. Specifically, all observability tools (1)  collect telemetry from the system (e.g., logs generated by an HTTP server, health-status of nodes over time), (2)  store them, allowing for queries on the data, and (3) generate alerts based on anomaly detection or defined rules (e.g., an excessive number of 404 HTTP errors within a time window).
2.2 Agentic AI
An LLM-based agent (or simply “agent”, hereafter) pairs an instruction-tuned LLM with a framework for autonomous interaction within a designed environment [54], enabling it to achieve objectives by planning, executing actions, and refining its strategy based on environmental feedback. This process uses pre-configured tools that the agent can call and configure to interact with the environment. Collectively, these capabilities form the agent’s action space.
While multiple agentic frameworks exist with varying levels of complexity, any agent can be abstracted as operating an iterative loop following three main steps. Provided with an task, an agent will act by:
  1. 1.
    Reasoning and Planning: The agent assesses the current state of the environment and designs the next actions, based on the currently available information.
  2. 2.
    Execution: The agent carries out the planned actions, which might result in a perturbation of the environment, e.g., performing a certain query on an observability tool or implement a firewall rule.
  3. 3.
    Responding: The agent considers the outcomes generated from the performed actions and updates its current information and beliefs.
This loop continues until an exit condition is reached, such as solving the given task or exhausting allocated resources (e.g., a set number of iterations or a time limit).
@@ 2.3 AIOps (AI for IT Operations)
Modern AIOps frameworks [11, 43, 53, 10, 52, 28, 50, 51, 47, 49, 45] are increasingly implemented using LLM-based agents. These agents gather and analyze telemetry from diverse sources, including system logs, performance metrics, traces, and alerts, to identify patterns, detect anomalies, and either suggest or carry out proactive and reactive actions. The overarching goal is to reduce downtime, improve response time, and lower operational costs compared to traditional human-driven approaches. Hereafter, we use the term “AIOps agent” to refer to an agent that performs an AIOps task.
AIOps typically operates in two main modes: (1) Human-in-the-loop, where the AI agent assists a human operator by generating analysis or recommendations, while a human (e.g., an on-call engineer) is responsible for executing remediation actions; and (2) Fully autonomous, where the AI agent handles the entire task automatically, without human intervention, in an end-to-end manner. Our study applies to both scenarios.
AIOps implementations can encompass a range of functionalities. In this work, we focus on the most common scenario: automated Root Cause Analysis (RCA) and incident response. We emphasize that RCA is a foundational component for any AIOps operations. Thus, our study directly impacts those operations, and AIOps as a whole.
@@ 2.3.1 Root Cause Analysis and Incident Response
Root cause analysis (RCA) is a structured, data-driven methodology aimed at identifying the fundamental causes of malfunctions, software bugs, or performance issues within IT systems. It involves systematically analyzing telemetry to trace problems to their origin. Once the root cause of an incident is detected, it is used to guide the response process; that is, implementing solutions to address the underlying malfunction (remediation).
In modern AIOps, RCA and remediation are implemented and automated by relying on AI agents. Provided with incident data, the agent is instructed to resolve the task using a set of diagnostic tools. Typically, its action space is defined by tailored function calls dedicated to streamlining telemetry collection from the observability stack running within the system (see Section 2.1), as well as access to general-purpose tooling such as direct shell access in order to get system-wide information and perform actions, e.g., checking firewall rules. An example of the agent’s execution is shown in Figure 1.
@@
    Activation: The agent is activated to perform the RCA task, typically in response to an alert or any signal that indicates a potential issue or anomaly in the system. The source of the alert can vary depending on the system’s implementation. Alerts may be automatically generated by observability tools based on predefined rules (e.g., a high number of password resets), or triggered by anomaly detection systems. In other cases, alerts may come directly from ticketing systems or be submitted through chat interfaces, such as a slack bot integrated with the agent [57, 42].
@@
    Analysis: Once an alert is received, it serves as the initial input for the agent. The agent begins its execution loop (see Section 2.2). This mainly involves querying the observability stack to collect telemetry associated with the alert and to query system information dynamically across multiple rounds (see Figure 1 for an example of execution).
@@
    Solution Submission: After gathering sufficient information about the incident’s origin, and once confident in its diagnosis and potential remediation, the agent proceeds to report its findings. This output can either be delivered to a human operator (e.g., an on-call engineer) for manual intervention or passed to another AIOps agent capable of carrying out automated remediation via a shell interface on a target machine.
Hereafter, we use the term AIOps agent to refer to an LLM agent that performs an AIOps task. Several AIOps agents have been proposed in both academia [11, 43, 53, 10, 52, 28, 50, 51, 47, 49, 45] and industry [13, 44, 19, 34, 42, 20, 32, 40, 57, 24, 18]. These implementations differ based on the agentic framework employed (e.g., ReAct [54] vs Flash [58]), the inclusion of additional modules such as memory or retrieval-augmented generation (RAG), the underlying LLMs used (e.g., GPT-4.1, Claude Sonnet 4), and the set of external tools accessible to the agent.
@@ 2.4 Prompt Injection
With direct prompt injection, an attacker directly feeds the LLM with manipulated input through interfaces like chatbots or APIs, with the goal generally being to misuse LLMs [30].
In contrast, indirect prompt injection targets external resources–such as web pages or databases–that the LLM accesses as part of its input processing, most frequently in retreival-augmented generation setups. Crucially, the external sources are often accessible to untrusted users, allowing attackers to indirectly plant malicious content. Such attacks have been shown to be effective in manipulating search systems [27, 46], disseminate propaganda [22, 55], various cybercrime strategies [22], or even used as defense against automated cyberattacks [37]. Further, unintended attacks have surfaced in production LLM-assisted search results [41], demonstrating how consequential these attacks can be.
@@ 2.5 Log Injection
Log injection[36] is a general term used to refer to vulnerabilities that arise when systems record untrusted input in logs without proper sanitization or encoding. This flaw can be exploited by attackers to alter the structure or content of application telemetry, e.g., log forging or log truncation[8]. The primary goal of such attacks is to manipulate the integrity of log data, often to conceal malicious activity by injecting misleading or disruptive entries. This can undermine incident response, corrupt audit trails, and hinder forensic investigations by making it difficult to distinguish legitimate events from falsified or malformed log records.
@@
Furthermore, our attacks technique extends beyond logs to include manipulation of other forms of telemetry, such as traces and metrics. To reflect this broader scope, we refer to our approach as telemetry injection.
@@ 3 AIOps as an Attack Vector
header.referrer = \$PAYLOAD
@@
to solve the compatibility problem..
@@
Figure 2: Stages of the proposed attack. In this example, the adversary’s remediation involves installing a vulnerable version of software on the system. Red components illustrate the flow of untrusted inputs (tainted telemetry effect) throughout the system.
@@ 3.1 Threat Model
Extending to Hardened AIOps. To validate that the proposed attacks are hard to mitigate using existing defense mechanisms, we also consider the scenario in which AIOps proactively deploys existing mechanisms to detect and mitigate attacks such as prompt injection. Specifically, Section 4.3 considers AIOps that are hardened with PromptShields, Prompt-Guard2 [6] [33], and DataSentinel [31].
Attacker’s Knowledge. $\mathcal{A}$ has no prior knowledge of the internal workings of $\hat{t}$ and does not know have specific information about the AIOps agent in use. This includes knowledge of the backend LLM(s) deployed by the system, the configuration or behavior of the underlying AIOps solutions, and which exact external inputs are incorporated into the system’s telemetry data.
Any insights the attacker gains about the target application are obtained either during the attack or through an initial reconnaissance phase. This may include probing the public interface of $\hat{t}$ (e.g., fuzzing, port scanning) to determine ($i$) which actions are permitted, ($ii$) what types of events are logged, (iiiiii) what anomalous behaviors may trigger alerts within the target.
Attacker’s Objective. Given access to the public interface of the target system (and with no knowledge of $\hat{t}$’s or AIOps internals), the attacker aims to drive $\hat{t}$ into an insecure state by exploiting the AIOps pipeline $\hat{t}$ relies on.
Specifically, the attacker’s strategy involves manipulating $\hat{t}$’s system state via legitimate actions, with the goal of influencing the AIOps agent to select a remediation action of the attacker’s choosing. This remediation is maliciously crafted to weaken the system’s security; e.g., by triggering the installation of a software version known to contain a remote code execution vulnerability, thereby enabling direct exploitation.
This state manipulation is carried out through a sequence of valid actions within the target system, such as issuing HTTP requests to specific URLs or invoking API calls, and requires no additional assumptions from $\mathcal{A}$.

### v2

@@ Abstract
In this work, we perform the first security analysis of AIOps solutions, showing that, once again, AI-driven automation comes with a profound security cost. We demonstrate that adversaries can manipulate system telemetry to mislead AIOps agents into taking actions that compromise the integrity of the infrastructure they manage. We introduce techniques to reliably inject telemetry data using error-inducing requests that influence agent behavior through a form of adversarial input we call adversarial reward-hacking–plausible but incorrect system error interpretations that steer the agent’s decision-making. Our attack methodology, AIOpsDoom, is fully automated–combining reconnaissance, fuzzing, and LLM-driven adversarial input generation–and operates without any prior knowledge of the target system.
@@ 1 Introduction
The increasing sophistication of software systems has led organizations to integrate AI deeply into IT operations [14, 44, 20, 34, 42, 21, 32, 40, 57, 25, 19], a paradigm widely known as AIOps, or AI for IT Operations [13]. At its core, AIOps leverages machine learning methods to automate tasks traditionally performed by human operators, including anomaly detection, incident diagnosis, and automated remediation [56]. The new wave of AIOps incorporates autonomous agents built upon large language models (LLMs), which dynamically interact with systems to diagnose issues and execute corrective actions. This rapid adoption of AI-driven automation promises significant benefits for IT Operations, notably reduced response times, increased efficiency, and lower operational costs; a vision shared by both academia [11, 43, 53, 10, 52, 28, 50, 51, 47, 49, 45] and industry [14, 44, 20, 34, 42, 21, 32, 40, 57, 25, 19].
@@
(In)security Through Automation: However, existing research [11, 43, 53, 10, 52, 28, 50, 51, 47, 49, 45] has generally overlooked the security impact that this automation introduces. AIOps agents rely heavily on the telemetry data, which they consume to make decisions. So far in this area, the telemetry data has been assumed to be faithful and trustworthy. In this work, we challenge the above assumption. We demonstrate that adversaries can reliably manipulate telemetry data to indirectly influence and ultimately control the behavior of AIOps agents. Specifically, we show that even in the most realistic threat models, in which the attacker knows nothing about the target infrastructure and holds no privileged access, they can pollute telemetry to bias AIOps systems into executing harmful remediations and compromising production systems. The consequences of this manipulation can be severe, enabling attackers to subvert otherwise secure systems through the very automation designed to serve them.
Tailored Injection Strategy: At a more technical level, we introduce AIOpsDoom; a practical and general automated attack framework designed to subvert AIOps systems. AIOpsDoom uses classical reconnaissance (e.g., port scanning and crawling) to collect information about the target solely through its publicly available interfaces. It then identifies which entry points are meaningful for injecting an adversarially crafted payload, so as to end up in the stored telemetry data that the AIOps agent will later process. Telemetry injection is achieved by intentionally performing malformed requests through fuzzing against the target system, in order to induce errors that are likely to be logged.
Custom Payloads: Following the information-gathering phase, AIOpsDoom automatically generates payloads specifically tailored to the target application, engineered to induce the agent into voluntarily transitioning the system into an insecure state. The payload design is inspired by the concept of reward hacking [7, 17], where an agent exploits underspecified objectives or environmental gaps to maximize rewards through low-effort yet reward-generating behaviors. Unlike classical prompt injection, which directly overrides an agent’s objectives, AIOpsDoom’s payloads subtly manipulate telemetry data, leading autonomous agents to independently select harmful actions under the belief that they are optimal solutions. This preserves the apparent legitimacy of the agent’s task (i.e., proposing or implementing a remediation for an incident) while covertly steering its behavior. For example, a seemingly benign log entry might suggest that a critical software downgrade is the appropriate fix for an anomaly, prompting the agent to autonomously trigger a damaging rollback. We evaluate our attack techniques against state-of-the-art models such as GPT-4o and GPT-4.1, and show that they can evade sophisticated prompt-defense solutions, including Microsoft’s PromptShields and Meta’s PromptGuard-2.
@@
  • We present the first security assessment of Agentic AI in the context of IT Operations, known as AIOps. Our end-to-end attack methodology integrates customized adversarial techniques, drawing inspiration from well-established attack vectors, software testing strategies, and reconnaissance principles. We implement and automate this methodology in AIOpsDoom, a tool that the community can use to evaluate the security of current AIOps solutions.
  • We develop a practical and general telemetry injection technique for reliably forging malicious telemetry in a target AIOps system, based on error-inducing requests and fuzzing.
@@
  • We propose AIOpsShield, a simple defense mechanism tailored for AIOps that sanitizes telemetry data by leveraging its structured nature and the minimal influence of user-generated content in legitimate incident management tasks. Empirical evaluations on established benchmarks show that AIOpsShield effectively blocks telemetry-based adversarial attacks without degrading agent performance.
@@ 2.1 Telemetry and Observability
Observability tools collect, analyze, and correlate telemetry data to provide insights into the internal states of IT systems. Any modern and sufficiently complex IT architecture today relies on some form of observability stack (a combination of multiple observability tools) to maintain a comprehensive view of the system, facilitate incident response, and enable effective root cause analysis. Collected data falls into three main categories:
@@
Numerous vendors offer observability solutions. While different products may feature unique functionalities, core capabilities typically remain consistent across these solutions. Specifically, all observability tools (1)  collect telemetry from the system (e.g., logs generated by an HTTP server, health-status of nodes over time), (2)  store them, allowing for queries on the data, and (3) generate alerts based on anomaly detection or defined rules (e.g., an excessive number of 404 HTTP errors within a time window).
@@ 2.2 AIOps (AI for IT Operations)
Modern AIOps frameworks [11, 43, 53, 10, 52, 28, 50, 51, 47, 49, 45] are increasingly implemented using LLM-based agents. These agents gather and analyze telemetry from diverse sources, including system logs, performance metrics, traces, and alerts, to identify patterns, detect anomalies, and either suggest or carry out proactive and reactive actions. The overarching goal is to reduce downtime and lower operational costs compared to traditional human-driven approaches. Hereafter, we use the term “AIOps agent” to refer to an agent that performs an AIOps task.
AIOps typically operates in two main modes: (1) Human-in-the-loop, where the AI agent assists a human operator by generating analysis or recommendations, while a human (e.g., an on-call engineer) is responsible for executing remediation actions; and (2) Fully autonomous, where the AI agent handles the entire task automatically, without human intervention, in an end-to-end manner. Hereafter, we abstract these two possibilities and design our attacks and defenses to be general and applicable to both scenarios (see Section 3.1).
@@ 2.2.1 Root Cause Analysis and Incident Response
Root cause analysis (RCA) is a structured, data-driven methodology aimed at identifying the fundamental causes of malfunctions, software bugs, or performance issues within IT systems; a key component of any AIOps task. It involves systematically analyzing telemetry to trace problems to their origin. Once the root cause of an incident is detected, it is used to guide the response process; that is, implementing solutions to address the underlying malfunction (remediation).
In modern AIOps, RCA and remediation are implemented and automated by relying on AI agents. Provided with incident data, the agent is instructed to resolve the task using a set of diagnostic tools. Typically, its action space is defined by tailored function calls dedicated to streamlining telemetry collection from the observability stack running within the system (see Section 2.1), as well as access to general-purpose tooling such as direct shell access in order to get system-wide information and perform actions, e.g., checking firewall rules. An example agent’s execution is shown in Figure 1.
@@
    Activation: The agent is activated to perform the RCA/incident response task, typically in response to an alert or any external signal that indicates a potential issue or anomaly in the system. The source of the alert can vary depending on the system’s implementation. Alerts may be automatically generated by observability tools based on predefined rules (e.g., a high number of password resets), or triggered by anomaly detection systems. In other cases, alerts may come directly from ticketing systems or be submitted through chat interfaces, such as a slack bot integrated with the agent [57, 42].
@@
    Analysis: Once an alert is received, it serves as the initial input for the agent. The agent begins its execution loop. This mainly involves querying the observability stack to collect telemetry associated with the alert and to query system information dynamically across multiple rounds (see Figure 1 for an example of execution).
@@
    Solution Submission: After gathering sufficient information about the incident’s origin, and once confident in its diagnosis and potential remediation, the agent proceeds to report its findings. This output can either be delivered to a human operator for manual intervention or passed to another AIOps agent capable of carrying out automated remediation via a shell interface on a target machine.
Several AIOps agents have been proposed in both academia [11, 43, 53, 10, 52, 28, 50, 51, 47, 49, 45] and industry [14, 44, 20, 34, 42, 21, 32, 40, 57, 25, 19]. These implementations differ based on the agentic framework employed (e.g., Flash [58]), the inclusion of additional modules such as memory or retrieval-augmented generation (RAG), the underlying LLMs used, and the set of external tools accessible to the agent.
@@ 2.3 Prompt Injection
Indirect prompt injection targets external resources–such as web pages or databases–that the LLM accesses as part of its input processing, most frequently in retrieval-augmented generation setups. Crucially, the external sources are often accessible to untrusted users, allowing attackers to indirectly plant malicious content. Such attacks have been shown to be effective in manipulating search systems [27, 46], disseminate propaganda [23, 55], various cybercrime strategies [23], or even used as defense against automated cyberattacks [37]. Further, unintended attacks have surfaced in production LLM-assisted search results [41], demonstrating how consequential these attacks can be.
@@ 2.4 Log Injection
Log injection[36] is a general term used to refer to vulnerabilities that arise when systems record untrusted input in logs without proper sanitization or encoding. This flaw can be exploited by attackers to alter the structure or content of application telemetry, e.g., log forging or log truncation[8]. The primary goal of such attacks is to manipulate the integrity of log data, often to conceal malicious activity by injecting misleading or disruptive entries. This can undermine incident response, corrupt audit trails, and hinder forensic investigations.
@@
Furthermore, our attack technique extends beyond logs to include manipulation of other forms of telemetry, such as traces and metrics. To reflect this broader scope, we refer to our approach as telemetry injection.
@@ 3 AIOps as an Attack Vector
header.referer = \$PAYLOAD
@@
to resolve the compatibility problem….
@@
Figure 2: Stages of the proposed attack. In this example, the adversary’s remediation involves installing a vulnerable version of software on the system (e.g., MySQL5.5). Red components illustrate the flow of untrusted inputs throughout the system.
@@ 3.1 Threat Model
Extending to Hardened AIOps. To validate that the proposed attacks are hard to mitigate using existing defense mechanisms, we also consider the scenario in which AIOps proactively deploys existing mechanisms to detect and mitigate attacks such as prompt injection. Specifically, Appendix D considers AIOps that are hardened with PromptShields [33], Prompt-Guard2 [6], and DataSentinel [31].
Attacker’s Knowledge. $\mathcal{A}$ has no prior knowledge of the internal workings of $\hat{t}$ and does not have specific information about the AIOps agent in use. This includes knowledge of the backend LLM(s) deployed by the system, the configuration or behavior of the underlying AIOps solutions, and which exact external inputs are incorporated into the system’s telemetry data. Furthermore, $\mathcal{A}$ ignores whether the target AIOps involves human-in-the-loop or is fully automated.
Any insights the attacker gains about the target application are obtained either during the attack or through an initial reconnaissance phase. This may include probing the public interface of $\hat{t}$ (e.g., fuzzing, port scanning) to determine ($i$) which actions are permitted, ($ii$) what types of events are likely to be logged, (iiiiii) what anomalous behaviors may trigger alerts within the target.
Attacker’s Objective. Given access to the public interface of the target system, the attacker aims to drive $\hat{t}$ into an insecure state by exploiting the AIOps pipeline $\hat{t}$ relies on. Specifically, the attacker’s strategy is to influence the target AIOps agent to select a malicious remediation action. This remediation is chosen to weaken the system’s security, for example by triggering the installation of a software version known to contain a remote code execution vulnerability, thereby enabling direct exploitation. To be effective against AIOps systems that involve human-in-the-loop, the malicious remediation is designed to appear realistic and resemble a common, valid solution.
Attacker’s Capabilities. The attacker has no privileged access to the target; the malicious remediation is induced by carrying out a sequence of valid actions within the target system, such as issuing HTTP requests to specific URLs or invoking API calls, and requires no additional assumptions from $\mathcal{A}$.